OpenStudy (anonymous):

Using 8 8's and addition only,can you make 1,000 ?

OpenStudy (anonymous):

888+8=896+8=904+8=912+88=1000

OpenStudy (anonymous):

Yeah that is right because i got 888+88+8+8+8

OpenStudy (anonymous):

Thanks @AidenS